OpenStudy (anonymous):

Solve the equation by extracting square roots. (4x + 7)2 = 52 Solve the equation by extracting square roots. (4x + 7)2 = 52 @Mathematics

OpenStudy (anonymous):

is it (4x+7)^2 or 2*(4x+7)?

OpenStudy (anonymous):

(4x + 7)^2 = 52

OpenStudy (anonymous):

16x^2+56x+49=52 16x^2+56x-3=0 use the formula to get x

jimthompson5910 (jim_thompson5910):

\[\Large (4x + 7)^2 = 52\] \[\Large 4x + 7 = \pm\sqrt{52}\] \[\Large 4x + 7 = \pm2\sqrt{13}\] \[\Large 4x = -7\pm2\sqrt{13}\] \[\Large x = \frac{-7\pm2\sqrt{13}}{4}\] So the solutions are \[\Large x = \frac{-7+2\sqrt{13}}{4} \ \text{or} \ x = \frac{-7-2\sqrt{13}}{4} \]